OpenStudy (anonymous):

Sally’s employer pays for 38% of her annual health insurance premium of $4,350.00. Sally pays the remaining balance by having it deducted in equal amounts from her 26 paychecks throughout the calendar year. How much will Sally have deducted from each paycheck? a. $63.58 b. $103.73 c. $137.75 d. $207.46

OpenStudy (anonymous):

what's 38% of 4350 ?

OpenStudy (anonymous):

4350 * 0.38 = 1653 this is what the employer pays for the health insurance Remaining amount is 4350 - 1653 = 2697 this is the amount that sally pays for the remaining part of heer insurance She divides that over 26 paychecks, i.e. 2697/26 = ? this will be the amount deducted from each of her paychecks Do u understand how we worked it our or shall I explain it further? thnx